OpenStudy (anonymous):

What is the solution of the system of equations? Show all work. 3x+2y+z=7 5x+5y+4z=3 3x+2y+3z=1

OpenStudy (anonymous):

@JakeV8 , forget the other one.. I need help with this one.

OpenStudy (anonymous):

@satellite73 HELP ME!

OpenStudy (anonymous):

@JakeV8 I need you.!

OpenStudy (anonymous):

@A_Level_Student can you please help me?!

jimthompson5910 (jim_thompson5910):

what method(s) are you allowed to use? if any, which methods are you familiar with?

OpenStudy (anonymous):

Substitution graphinf elimination ect.. And I just need someone to take me through it step by step, I've been working on this homework for 6 hours....

jimthompson5910 (jim_thompson5910):

Substitution may be easiest...graphing is most likely impossible (unless you have a 3D graphing calculator with you)

jimthompson5910 (jim_thompson5910):

What do you get when you solve 3x+2y+z=7 for z ??

OpenStudy (anonymous):

idk... i kept getting like ridiculous negative decimals.....

jimthompson5910 (jim_thompson5910):

Subtract 3x from both sides 3x+2y+z=7 2y+z=7-3x Then subtract 2y from both sides 2y+z=7-3x z=7-3x-2y Then optionally, rearrange the terms z=7-3x-2y z=-3x-2y+7 So after solving 3x+2y+z=7 for z, we get z=-3x-2y+7

jimthompson5910 (jim_thompson5910):

With this, we can replace each 'z' in the other two equations with -3x-2y+7 to get Plug in z=-3x-2y+7 into the second equation 5x+5y+4z=3 5x+5y+4(-3x-2y+7)=3 5x+5y-12x-8y+28=3 -7x-3y+28=3 -7x-3y=3-28 -7x-3y=-25 So after replacing 'z' with -3x-2y+7 in the second equation 5x+5y+4z=3, we get -7x-3y=-25 ------------------------------------------------------- Then do the same for third equation 3x+2y+3z=1 3x+2y+3(-3x-2y+7)=1 3x+2y-9x-6y+21=1 -6x-4y+21=1 -6x-4y=1-21 -6x-4y=-20 So after replacing 'z' with -3x-2y+7 in the third equation 3x+2y+3z=1, we get -6x-4y=-20

jimthompson5910 (jim_thompson5910):

This is a lot, but we come away with the new system of equations -7x-3y=-25 -6x-4y=-20 You can use any method you want to solve this new system of equations to find x and y.

OpenStudy (narissa):

is it 4,-3